# Conceptualization, Methodology, Writing – original draft, Data curation, Software, Writing – review & editing

### v1

@@ 6.3 Sensitivity analysis
To this end, we analyze the system using both the linear method and the proposed IDF method, considering both state and control input saturation, with the saturation limits set as specified in Subsections 6.1 and 6.2, and using the default controller parameters. The frequency response is examined over a range from $0.1$ to $0,5Hz$ and an extended range of $R/a_{bound}$ from 00 to $8$ (correspondingly, 00 to $4$ for $R/v_{bound}$). The results are presented as heat maps in Fig. 6.3, where darker colors represent smaller magnitudes and more negative phases. From, Fig. 6.3(a)-(b), using the linear method, we observe that the magnitude and phase remain unchanged with varying $R/a_{bound}$ since the linear method does not account for saturation nonlinearity. However, as shown in Fig. 6.3 (c)-(d), given the state and control input limits, the frequency response is in fact dependent on the oscillation amplitude of the predecessor. Generally, when the saturation limits are reached, the magnitude decreases and the phase becomes more negative compared to the linear method. This indicates that the linear method overestimates the amplification and underestimates the response time for an AV when these saturation limits are present. As $R/a_{bound}$ increases, the frequency at which the saturation limits are reached decreases. Furthermore, for the same value of $R/a_{bound}$, as the frequency increases, the magnitude gradually decreases after the limits are reached, while the phase undergoes a sudden shift when $R/a_{bound}$ exceeds approximately 5. These patterns are consistent with and generalize the results illustrated in Subsections 6.1 and 6.2.
@@
For clearer visualization of the discrepancies, the magnitude difference, $|F|_{lin}-|F|_{IDF}$, and phase difference, $\angle F_{lin}-\angle F_{IDF}$, are shown in Fig. 6.3. Here, the subscripts ”$lin$” and ”$IDF$” represent results obtained using the linear and $IDF$ methods, respectively. The regions where the limits are not reached and where they are reached are separated by a red line and appropriately labeled.
@@ 7 Conclusion
The proposed method is evaluated for AV systems integrated with solely the control saturation, solely the state saturation, and a combination of both. The theoretical frequency response results are consistent with the Simulink estimations and surpass other theoretical frameworks in terms of analyzing the oscillation attenuation and response time of AVs. It is found that when a saturation limit is reached, generally speaking, the widely used linear string stability analysis tends to underestimate both the attenuation effect and response time of AVs, when both traffic state and control input limits are reached, this phenomenon is even more pronounced. Our study also reveals that when saturation limits are reached for all frequencies, the commonly used linear method can misjudge string stability, while the proposed method remains accurate. Future work includes investigating highly asymmetric limits using the dual-input describing function.

### v3

@@ 6.3 Sensitivity analysis
To this end, we analyze the system using both the linear method and the proposed IDF method, considering both state and control input saturation, with the saturation limits set as specified in Subsections 6.1 and 6.2, and using the default controller parameters. The frequency response is examined over a range from $0.1$ to $0,5Hz$ and an extended range of $R/a_{bound}$ from 0 to $8$ (correspondingly, 0 to $4$ for $R/v_{bound}$). The results are presented as heat maps in Fig. 6.3, where darker colors represent smaller magnitudes and more negative phases. From, Fig. 6.3(a)-(b), using the linear method, we observe that the magnitude and phase remain unchanged with varying $R/a_{bound}$ since the linear method does not account for saturation nonlinearity. However, as shown in Fig. 6.3 (c)-(d), given the state and control input limits, the frequency response is in fact dependent on the oscillation amplitude of the predecessor. Generally, when the saturation limits are reached, the magnitude decreases and the phase becomes more negative compared to the linear method. This indicates that the linear method overestimates the amplification and underestimates the response time for an AV when these saturation limits are present. As $R/a_{bound}$ increases, the frequency at which the saturation limits are reached decreases. Furthermore, for the same value of $R/a_{bound}$, as the frequency increases, the magnitude gradually decreases after the limits are reached, while the phase undergoes a sudden shift when $R/a_{bound}$ exceeds approximately 5. These patterns are consistent with and generalize the results illustrated in Subsections 6.1 and 6.2.
@@
For clearer visualization of the discrepancies, the magnitude difference, $|F|_{lin}-|F|_{IDF}$, and phase difference, $\angle F_{lin}-\angle F_{IDF}$, are shown in Fig. 6.3. Here, the subscripts ”linlin” and ”IDFIDF” represent results obtained using the linear and $IDF$ methods, respectively. The regions where the limits are not reached and where they are reached are separated by a red line and appropriately labeled.
@@ 7 Conclusion
The proposed method is evaluated for AV systems integrated with solely the control saturation, solely the state saturation, and a combination of both. The theoretical frequency response results are consistent with the Simulink estimations and surpass other theoretical frameworks in terms of analyzing the oscillation attenuation and response time of AVs. It is found that when a saturation limit is reached, generally speaking, the widely used linear string stability analysis tends to underestimate both the attenuation effect and response time of AVs. When both traffic state and control input limits are reached, this phenomenon is even more pronounced. Our study also reveals that when saturation limits are reached for all frequencies, the commonly used linear method can misjudge string stability, while the proposed method remains accurate. Future work includes investigating highly asymmetric limits using the dual-input describing function.